These are the Actions
How many ways are there of selecting one of each face card from the pile?
How many ways are there of selecting the queen of clubs, then the king of diamonds, and then the jack of hearts from the pile?
How many ways are there of selecting three of the same face cards (i.e., 3 jacks, 3 queens, or 3 kings) from the pile?
This is what I need to know Derivation and solution Combination or permutation Why
Thanks for you help

Answer by stanbon(75887)

These are the Actions
How many ways are there of selecting one of each face card from the pile?
There are three types of face card, K, Q, J.
There are four of each type.
#of ways to choose a "K" is 4C1 = 4
#of ways to choose a "Q" is 4C1 = 4
#of ways to choose a "J" is 4C1 = 4
# of ways to choose one of each is 4*4*4= 64
----------------
How many ways are there of selecting the queen of clubs, then the king of diamonds, and then the jack of hearts from the pile?
One way to choose queen of clubs
One way to choose king of diamonds
one way to choose jack of hearts
# of ways to do this, and this, and this is 1*1*1=1
Note: order makes no difference.
---------------------
How many ways are there of selecting three of the same face cards (i.e., 3 jacks, 3 queens, or 3 kings) from the pile?
1st: select one of the three types = 3 ways
2nd: select 3 of the 4 cards of that type = 4C3=4
Answer to your problem: 3*4=12 ways
----------------------------
Cheers,
Stan H.